Morton Company’s contribution format income statement for last month is given below:

Sales (15,000 units × $30 per unit) $450,000

Variable expenses 315,000

Contribution margin 135,000

Fixed expenses 90,000

Net operating income$45,000

1. New equipment has come onto the market that would allow Morton Company to automate a portion of its operations. Variable expenses would be reduced by $9 per unit. However, fixed expenses would increase to a total of $225,000 each month. Prepare two contribution format income statements, one showing present operations and one showing how operations would appear if the new equipment is purchased.

2. Refer to the income statements in (1). For the present operations and the proposed new operations, compute (a) the degree of operating leverage, (b) the break-even point in dollar sales, and (c) the margin of safety in dollars and the margin of safety percentage.

3. Refer again to the data in (1). As a manager, what factor would be paramount in your mind in deciding whether to purchase the new equipment? (Assume that enough funds are available to make the purchase.)

4. Refer to the original data. Rather than purchase new equipment, the marketing manager argues that the company’s marketing strategy should be changed. Rather than pay sales commissions, which are currently included in variable expenses, the company would pay salespersons fixed salaries and would invest heavily in advertising. The marketing manager claims this new approach would increase unit sales by 30% without any change in selling price; the company’s new monthly fixed expenses would be $180,000; and its net operating income would increase by 20%. Compute the company's break-even point in dollar sales under the new marketing strategy.
Answer #1

1.Contribution Income Statement

A Contribution income statement is an income statement in which all the variable expenses are deducted from sales to arrive at a contribution margin, from which all fixed expenses are subtracted to arrive at the net profit or net loss for the period.

New Proposed Operations ( purchase of new equipment)

Number of units sold= 15000 units

Variable Expenses will be reduced $9 per unit= 15000*9=$135000

Fixed Expenses = $ 225000


Amount in $

Contribution Income Statement for the Month Present Proposed(new)

Sales (15000 units * $30 ) 4,50,000 4,50,000

Less Variable expenses 3,15,000 1,35,000

Contribution Margin 1,35,000 3,15,000

Less Fixed expenses 90,000 2,25,000

Net Operating Income 45,000 90,000

2

a.Degree of Operating leverage(DOL)

     The degree of Operating leverage is the leverage ratio that sums up the effect of an amount of operating leverage on the company's Earning Befor Interest and tax.(EBIT)

DOL= Contribution Margin  / Operating Income

Present Operations

DOL = 135000/45000=3times

Proposed Operations

DOL=315000/90000=3.5 times

b.Break even Sales

Break even sales is the dollar amount of revenue at which a business earns a profit of zero.

Break even sales= Fixed Expenses / Contribution Margin Percentage

Contribution margin percentage = Contribution / sales *100%

Present Operations

Contribution margin percentage = 135000/450000 *100=30%

Break even sales = 90000/30%=$300,000

  Proposed Operations

Contribution margin percentage = 315000/450000*100=70%

Break even sales = 225000/70%=$321,428.57

c.Margin of Safety in Dollars and Percentage

Margin of safety is built in cushion allowing for some losses to be incurred without major negative effect.

Margin of Safety in Dollars = Current Sales - Break Even Sales

Margin of Safety in percentage=(Current Sales - Break Even Sales) /Current Sales *100

Present Operations

  Margin of Safety in Dollars =450000-300000=$150000

Margin of Safety in percentage=150000/450000*100=33.33 %

Proposed Operations

  Margin of Safety in Dollars =450000-321428.57=$128,571.43

Margin of Safety in percentage=128571.43/450000*100=28.57%

3.Factor that paramount in deciding whether to purchase the new equipment.(other than fund)

The important factor in deciding whether to purchase a new equipment is the cost-benefit anlysis of that purchase. It means if we are purchasing it ,then there should be increase in revenue and additional costs incurred should be compensated by increasing sales. The other factors which influence the purchasing decision are as follows,

New or used machine should be puchased.

The Impact of purchase in production output.

The price Comparison

Manpower needed for the new equipment.

Space and power requirement of the new equipment.

Installation and warranty period of the new equipment.

4.Break even sales in dollars of the new proposed marketing strategy

Increase in sales = 30% increase

= 450000+ 30% =$585,000

New Fixed expenses= $180,000

New operating income= 20% increase

= 45000+20%=$54,000

Break even sales= Fixed Expenses / Contribution Margin Percentage

Contribution = Fixed expense + Operating income

= 180000+54000=$234,000

Contribution margin percentage= Contribution/sales*100

= 234000/585000*100=40%

Break Even sales in dollars = 180000/40%=$450,000